OpenStudy (anonymous):

Write the equation in slope intercept form of the line that is parallel to y=-3x+5 and contains (2,-3)?

OpenStudy (anonymous):

i WILL GIVE ANYONE A MEDAL WHO HELPS ME WITH THIS QUESTION

OpenStudy (anonymous):

First of all, if it is parallel to y = -3x + 5, then it has the same slope.

OpenStudy (anonymous):

So, we have y = -3 x + b, and it goes through the point (2, -3) so we plug the x and y values into to the equation to find b.

OpenStudy (anonymous):

-3 = -3 * 2 + b -3 = -6 + b ( add 6 to both sides) b = 3 so the equation is y = -3x + 3

OpenStudy (anonymous):

thankyou, can you answer another question similar to the one you just answered?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

Write the equation in slope intercept form for the line that contains (-4,-2) and is perpendicular to the line y=1/4x-7

OpenStudy (anonymous):

If the line is perpendicular, the the slope is equal to the negative reciprocal of the line you have. y = 1/4 x - 7 In other words the slope for the line you are looking for is -4 y = -4 x + b The rest is the same, plug in the x and y values from your point and solve for b.

OpenStudy (anonymous):

Think you can handle it from there?

OpenStudy (anonymous):

is the Answer y= -4x-8 ?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

u wanna help me please? i have a test tommorrow?

OpenStudy (anonymous):

it's y = -4 x + 14 Can you show me how you got b = -8, I might be able to figure out where you're mistake was